Systems Staff Meeting Notes 1/26/98

Systems Staff Meeting Notes
January 26, 1998

Present: Rich Bennett, Bill Covey, Destry Fudge, Debra Harris, Winston Harris, Sharon Henry, Bobbie Parker, Suzy Shaw, Leota Whitfield

Announcements
  1. Bill recently talked with Tom Cetwinski. Seems the position procedures have been changed & simplified. Bobbie Parkerís position is being reclassified directly. Lonís old position will be reclassified to a Computer Programmer Analyst and it has been proposed that Brian Burton will fill that position. Should be all done by the end of the week - as long as money boundaries are maintained.
  2. Jason Monahanís last day is slated for February 4 (next Wednesday). Bill was recently over at MSL and was impressed with the computer knowledge of some of the students working over there. May try to recruit one of them as Jasonís replacement.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. IBM sent a CD for the Networking Work Station - it arrived last Tuesday afternoon. It is an Ďevaluation productí - not an official one, and came at no charge. Bill is trying to make it fly - the CD was marked as a 5th edition & requires a java script browser.
  2. CISCO came by to pick up the switch order - they had not received it yet and wanted to hand deliver it.
  3. Lee received a zip drive back from Jack Waters - he said it never worked properly. Lee notes it was purchased in 1996. He says it shuts down after 15 minutes. We will try it out here in Systems - will need to recoup the cartridges. It may just need cleaning, there isnít much to be done in the way of fixing them. Lee says it has a rattle to it.
  4. Destry heard from OED - told him to put a set file in the autoexec. Destry can make it work by adding ini.files to all Windows directory, throwing set file in & adding to WIN.iri. Bill says this is not an acceptable fix. Win.iri controls how the machine comes up. Current stuff is not the network version.
  5. IP 238 needs another range added to feed the new SWAMP area.
  6. MSL needs new subnet added.
Liaisons
  1. EMAIL address list is not back yet.
  2. Memory lists have been received from DOCS & ILL only.
Networking
  1. Campus Networking - still floundering with the idea of networked walk-up plug-ins. Students are currently walking up at MSL with orange extension cords in tow for their laptops. Money to fund this project has dissipated, but the Directors have requested that Bill approach Capaldi for funding for the Library plug-ins. HVAC & building power should not be a problem at our level. Will probably provide this service only in Room 100.
  2. The Circulation desk in Library West is soon to be renovated - probably during Spring Break. Currently, the desk is scheduled to be taken down on Friday, March 6 at 6 PM. They will keep regular hours on Saturday, then install the new desk on Sunday. All crates will be stored in the lobby prior to assembly. The gates will be closer to the desk, and will need to be put on a separate circuit. Bill says there are 3 lines coming down from the closet - weíll need to check the length. Will add Hub & jumpers and need to have one SNA up for at least one week after the desk installation.
  3. Marcia was looking for Welch Allen information - for Circ desk scanners - this is no longer the vendor, Lee referred her to Cataloging.
  4. Rich will need IPís for the Circ desk by mid-February to get the table changes in to FCLA to provide tech access.
  5. Work station allocations - 13 Cows have been not yet been assigned, if they are not to be used for this project, weíll have to get another order out.
  6. Special Collections - wiring of bridge & annex should be done by tomorrow (Tuesday).
  7. LAC, N. Tower - Wiring status unknown.
  8. Education - some errors discovered in the job: had the CAT5 connector at the top of a 20í ceiling, needs to be brought down to be able to reach it. Hopefully, job will be done this week.
  9. Systems & Preservation - Martha is wanting estimates, Destry is working on diagrams - then comes the bidding & purchasing process. The plan is to combine this project with the digital phone wiring system.
  10. MSL - awaiting estimate - PR electrical checking - it will be bad. The original drawings didnít define the work stations.
Software Projects Status
  1. ILL on hold - Winston is working on OPS tracking.
  2. Sharon is working on the OPS side of the Personnel migration.
  3. ATC standalone project on hold.
  4. Influx of app sharing - classroom will be next, using PowerPoint.
DP Coordinator
  1. Systems modular furniture is being worked on today - wall going up between us & Preservation.
  2. John Humphrey is getting estimates on the electrical wiring for the modular area.
  3. Digital phone estimates are in - thereís a question of combining digital & key system on the same switch.
Calendar
  1. Bill Covey out Jan. 29 & 30.
  2. Winston Harris out February 1.


Back to the Systems Homepage

Last updated March 7, 2000
by Debra Harris
debharr@mail.uflib.ufl.edu